The equivalent resistance between points A and B of the circuit shown is 25.2 Ω. For R1 = 10.3 Ω and R2 = 52.8 Ω, calculate the value of resistance R3.

A B 1
Answer #1

R2 and R3 are in parallel

net resisrance=R2*R3/(R2+R3)

R1 is in series with the above equivalent resistance.

equivalent resistance between A and B =R1+(R2*R3/(R2+R3))

using the given values,

10.3+(52.8*R3/(52.8+R3))=25.2

==>52.8*R3/(52.8+R3)=14.9

==>3.54362*R3=52.8+R3

==>R3=52.8/(3.54362-1)=20.7578 ohms

hence value of resistance of R3 is 20.7578 ohms